# Quantum Gravity

### v1

@@ 2.4 Loop integrals in nonlocal theories
where the polynomial $P(p)$ is given by (2.2). Associated with $G_{\text{nl}}(p)$, we have a cone $\mathcal{C}$ where the expansion (2.7) applies.
Consider a Feynman diagram $F$. Let $p,p_{i}$ denote the internal momenta and $k,k_{a}$ the external ones. The loop integral associated with $F$ is the integral of a product of propagators $G_{\text{nl}}(p_{i})$ times a product of functions $V(p_{i},k_{a})$ originated by the vertices. The latter are under control, as we explain in subsection 6.4, so we concentrate on the propagators.
@@
It is easy to show that the integrals are convergent for Euclidean $k_{a}$, in the sense of the dimensional regularization. Consider, for example, the expression (2.23) for $\mathcal{B}(k)$. For Euclidean $p$ and $k$, the arguments $P(p)$ and $P(p+k)$ are located inside the cones $\mathcal{C}$ and $\mathcal{C}^{\prime}$ associated with the propagators $G_{\text{nl}}(p)$ and $G_{\text{nl}}(p+k)$. Formula (2.7) then ensures that (2.23) tends to the same integral with $G_{\text{nl}}(p)\rightarrow P(p)$ and $G_{\text{nl}}(p+k)\rightarrow P(p+k)$, which is indeed convergent in the sense of the dimensional regularization (assuming that the vertices do not invalidate the argument, see 6.4). What this means is that there exists an open set $\Omega$ of the complex $D$ plane where the integral is convergent in the standard sense, or can be split into a finite sum of integrals that admit convergence domains $\Omega_{i}$ in the standard sense [23].
Now we prove that the loop integrals are convergent, in the sense of the dimensional regularization, for every complex external momenta $k_{a}$. Actually, they are entire functions of $k_{a}$, so it is possible to replace the external momenta $k$ with their Minkowski versions $k_{\text{M}}$ directly inside the integrals.
@@
$P(p+k)=(p+k)^{2}+m^{2}=\rho\mathrm{e}^{i\theta},\qquad\theta=\arctan\frac{\text{Im}[k_{4}(k_{4}+2p_{4})]}{\text{Re}[(p+k)^{2}+m^{2}]}.$
The phase $\theta$ of $P(p+k)$ tends to zero when the integrated momentum $p$ tends to infinity, in any direction. This implies that the argument of $G_{\text{nl}}(p+k)$ falls off fast enough inside $\mathcal{C}^{\prime}$ for sufficiently large $p$, which makes the integral (2.24) convergent.
Together with the fact that the integrand is obviously regular everywhere, this argument proves that the function $\mathcal{B}(k)$ is entire. Then we can perform the analytic continuation E $\rightarrow$ M by replacing $k=(k_{4},\mathbf{k})$ with its Minkowskian version $k_{\text{M}}=(-ik^{0},\mathbf{k})$ directly inside. The Minkowskian bubble diagram is thus
@@
The property just proved extends to more complex loop integrals, which also define entire functions of $k_{a}$. Note that it does not apply, on the contrary, to local quantum field theory. There, the entire functions $G_{\text{nl}}$ are replaced by the reciprocals of polynomials. The integrands have poles, which can cross the integration path when $k$ is deformed to complex values. Direct replacements $k\rightarrow k_{\text{M}}$ jump over the integration paths, giving incorrect results.
@@ 3 Local limit in Euclidean space
At the tree level, the assumption (2.4) ensures that the local limit of the Lagrangian (2.1) is the Lagrangian of a local theory:
@@
In what follows, we assume, again, that the vertices are local or their limits as $\lambda\rightarrow\infty$ are smooth enough to not affect our arguments. The theories we have in mind satisfy this assumption, as we demonstrate in subsection 6.4.
@@
It is also straightforward to show that the correlation functions of the nonlocal theory tend to those of the local theory (3.1). Indeed, the loop integrals just involve the Euclidean domain. There, the values of the rescaled polynomial $P$ are always real and larger than a positive number, $\lambda P(0)=\lambda m^{2}$, so the expansion (2.7) can be used to prove the statement.
@@
Since $k$ is Euclidean, both $P(p)$ and $P(p-k)$ have Euclidean arguments.
@@
is convergent in the physical dimension $D=4$. The bounds (2.11) and (2.16) allow us to apply the dominated convergence theorem and conclude that the limit (3.3) gives $\mathcal{B}_{\text{L}}(k)$.
@@
[…]
@@ 4 Local limit in Minkowski spacetime
The next task is to investigate the local limit of the Minkowskian correlation functions. We cannot just compose the map (3.5) with the continuation E $\rightarrow$ M, because the continuation of a local theory admits a plurality of choices:
@@
The simplest option is the analytic continuation, which is excluded by unitarity, as we show below in detail. Among the other options, a special place is reserved to the “average continuation”, which defines purely virtual particles [11]. The worst possibility is that the local limit in Minkowski spacetime does not even exist. In particular, we cannot expect that the limit of (2.24) is just (3.4) with $k\rightarrow k_{\text{M}}$, which is ill defined.
@@ E $\downarrow$ M E $\downarrow$ M
Assume for the moment that the vertices $V_{1}$ and $V_{2}$ are identically one. Making a reflection $p_{4}\rightarrow-p_{4}$ in (2.24) we infer that $B_{\text{M}}(k_{\text{M}})=B_{\text{M}}(k_{\text{M}}^{\ast})$. Because of the property (2.10) and the overall factor $i$, the conjugate $B_{\text{M}}^{\ast}(k_{\text{M}})$ coincides with $-B_{\text{M}}(k_{\text{M}}^{\ast})$. Hence, $B_{\text{M}}(k_{\text{M}})=-[B_{\text{M}}(k_{\text{M}})]^{\ast}$. Given that the real part of $B_{\text{M}}(k_{\text{M}})$ vanishes identically, its local limit must vanish as well, if it exists.
Thus, the local limit of $B_{\text{M}}(k_{\text{M}})$ is purely imaginary. If it contained propagating particles or ghosts, its real part would be nontrivial. On the contrary, it is allowed to contain PVPs, because they give a purely imaginary bubble diagram (check the appendix).
@@
A more general version of the argument is based on the optical theorem, which reads $-iT+iT^{\dagger}=TT^{\dagger}$, where $S=1+iT$ is the $S$ matrix and $iT$ collects the loop diagrams in matrix form. Consider the “empty” theory (2.1). We call it this way, because it does not propagate any degree of freedom, differently from (2.3). This means $TT^{\dagger}=0$, hence the matrix $T$ is Hermitian: all the loop diagrams are purely imaginary.
Since Re[$iT$]$=0$ for arbitrary finite $\lambda$, the local limit $\lambda\rightarrow\infty$ can only return a local theory with Re[$iT$]$=0$, that is to say, still an empty theory. No physical particles or ghosts can be generated by the limit. Purely virtual particles are allowed, precisely because they satisfy Re[$iT$]$=0$.
@@
$\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}=\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}+\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\hskip 0.72229pt.$ (4.3)
@@
$0=\text{Re}\left[\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}\right]=\text{Re}\left[\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}\right]+\text{Re}\left[\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\right].$ (4.4)
Assume, ad absurdum, that the local limit adds degrees of freedom. Then the optical theorem $-iT+iT^{\dagger}=TT^{\dagger}=-2\text{Re}$[$iT$], applied to the limit itself, tells us that $TT^{\dagger}$ is nonzero. For example, in the case of the ordinary bubble diagram with circulating physical particles, formula (A.1) gives
[…]
@@
[…]
@@
$\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}=\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{PVP}}+\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}},$ (4.5)
and recalling that the bubble diagram (A.2) with circulating PVPs is purely imaginary, the real part of the rest vanishes,
$0=\text{Re}\left[\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}\right]=\text{Re}\left[\hskip 0.72229pt\rangle{\bigcirc}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\right],$
@@
We stress once again that the results of this section are predicated on the assumption that the local limit exists. It turns out that it does with the second choice of (2.5) for the function $Q(P)$, while it does not with the first choice.
5 Local limit of Minkowskian correlation functions
In this section we study the local limit in Minkowski spacetime with the options (2.5). We start from the tree level, then proceed to analyze the bubble and triangle diagrams in detail. We show that the local limit is singular with the left option and tends to a local theory with PVPs with the right option, albeit with some caveats.
[…]
$G_{\text{nl}}^{\text{M}}(p_{\text{M}})=-\frac{i\tau}{Q(P(p_{\text{M}}))}.$ (5.1)
At the tree level, formulas (2.8) and (2.17) give
$\lim_{\lambda\rightarrow+\infty}-\frac{i\tau\lambda}{Q(\lambda P(p_{\text{M}}))}=\left\{\begin{tabular}[]{l}$-\frac{i\tau}{|P(p_{\text{M}})|}$\\ $-\mathcal{P}\frac{i\tau}{P(p_{\text{M}})}$\end{tabular}\right.\begin{tabular}[]{l}for $Q(P)=h(P),$\\ for $Q(P)=\frac{h^{2}(P)}{P}.$\end{tabular}$ (5.2)
The top result illustrates the problem with the left choice of (2.5): the $\lambda\rightarrow+\infty$ limit does not give an acceptable propagator for a local theory. The second choice for (2.5), on the other hand, gives the answer we expect for a PVP in the classical limit.
It is important to stress that the principal value is not the right answer for PVPs at the level of radiative corrections. If we adopt it as an alternative propagator in Feynman diagrams, we obtain a theory of “Wheelerons” [24], which propagates ghosts. For example, the bubble diagram ends up giving the integral
$\mathcal{P}\int\frac{\mathrm{d}^{D}p_{\text{M}}}{(2\pi)^{D}}\frac{1}{p_{\text{M}}^{2}-m^{2}}\frac{1}{(p_{\text{M}}-k_{\text{M}})^{2}-m^{2}},$ (5.3)
which has a nonvanishing, unphysical absorptive part [25], absent in (A.2).
In a theory of PVPs, the radiative corrections are not given by the usual diagrams with a different propagator, but by new combinations of diagrams [10]. In view of this, before concluding that the local limit with the second choice of (2.5) gives a theory with PVPs, we must prove that the loop diagrams somehow “know what to do” by themselves. That is to say, for some mysterious reason they are already equipped with the instructions to implement the diagrammatic rules of [11, 12, 10], instead of those that lead to (5.3). Luckily, this turns out to be true [^4] Besides, we already know that they cannot give (5.3), because unitarity rules out local limits with ghosts..
5.1 Bubble diagram

### v3

@@ 2.2 Propagator
where the polynomial $P(p)$ is given by (2.2), the function $Q$ is given in (2.6) and $h$ given by (A.1). In the local limit, both choices (2.6) give $\tau/P$ in Euclidean space. As we are going to show, the two options lead to crucially different results in Minkowski spacetime.
Associated with $G_{\text{nl}}(p)$, we have a double cone $\mathcal{C}$ where the expansion (A.2) applies.
2.3 Loop integrals
Now we study general properties of the Feynman diagrams of nonlocal theories. Let $p,p_{i}$ denote the internal momenta and $k,k_{a}$ the external ones. The loop integral associated with a diagram $F$ is the integral of a product of propagators $G_{\text{nl}}(p_{i})$ times a product of functions $V(p_{i},k_{a})$ originated by the vertices. The latter are also entire functions, as we explain in subsection 5.4, so we concentrate on the propagators.
@@ 2.3 Loop integrals
It is easy to show that the integrals are convergent for Euclidean $k_{a}$, in the sense of the dimensional regularization. Consider, for example, the expression (2.8) for $\mathcal{B}(k)$. For Euclidean $p$ and $k$, the arguments $P(p)$ and P(pk)P(p-k) are located inside the cones $\mathcal{C}$ and $\mathcal{C}^{\prime}$ associated with the propagators $G_{\text{nl}}(p)$ and $G_{\text{nl}}(p-k)$. Formula (A.2) then ensures that (2.8) tends to the same integral with $G_{\text{nl}}(p)\rightarrow P(p)$ and $G_{\text{nl}}(p-k)\rightarrow P(p-k)$, which is indeed convergent in the sense of the dimensional regularization (assuming that the vertices do not invalidate the argument, see 5.4). What this means is that there exists an open set $\Omega$ of the complex $D$ plane where the integral is convergent in the standard sense, or that the integral can be split into a finite sum of integrals that admit convergence domains $\Omega_{i}$ in the standard sense [22].
Now we prove that the loop integrals are convergent, again in the sense of the dimensional regularization, for every complex external momenta $k_{a}$. Actually, they are entire functions of $k_{a}$, so it is possible to replace the external momenta $k$ with their Minkowski versions $k_{\text{M}}$ directly inside the integrals.
@@
$P(p-k)=(p-k)^{2}+m^{2}=\rho\mathrm{e}^{i\theta},\qquad\theta=\arctan\frac{\text{Im}[k_{4}(k_{4}-2p_{4})]}{\text{Re}[(p-k)^{2}+m^{2}]}.$
The phase $\theta$ of P(pk)P(p-k) tends to zero when the integrated momentum $p$ tends to infinity, in any direction. This implies that the argument of $G_{\text{nl}}(p-k)$ falls off fast enough inside $\mathcal{C}^{\prime}$ for sufficiently large $p$, which makes the integral (2.9) convergent.
Together with the fact that the integrand is regular everywhere (both the vertices and the propagators being entire functions, see section 5.4), this argument proves that the function $\mathcal{B}(k)$ is entire. Then we can perform the analytic continuation E $\rightarrow$ M by replacing $k=(k_{4},\mathbf{k})$ with its Minkowskian version $k_{\text{M}}=(-ik^{0},\mathbf{k})$ directly inside. The Minkowskian bubble diagram is thus
@@
The property just proved extends to more complex loop integrals, which also define entire functions of $k_{a}$. Note that it does not apply, on the contrary, to local quantum field theory. There, the entire functions $G_{\text{nl}}$ are replaced by the reciprocals of polynomials. The integrands have poles, which can cross the integration path when $k$ is deformed to complex values. Direct replacements $k\rightarrow k_{\text{M}}$ may jump over the integration paths, and give incorrect results.
@@ 3 Local limit in Euclidean space
At the tree level, the assumption (2.5) ensures that the limit of the Lagrangian (2.1) is the Lagrangian of a local theory:
@@
We assume that the vertices tend to local limits smoothly enough to not affect our arguments. The theories we have in mind satisfy this assumption, as we demonstrate in subsection 5.4.
@@
It is also straightforward to show that the correlation functions of the nonlocal theory tend to those of the local theory (3.1). Indeed, the loop integrals just involve the Euclidean domain. There, the values of the rescaled polynomial $P$ are always real and larger than a positive number, $\lambda P(0)=\lambda m^{2}$, so the expansion (A.2) can be used to prove the statement.
@@
Since $k$ is Euclidean, both $P(p)$ and P(pk)P(p-k) have Euclidean arguments.
@@
is convergent in the physical dimension $D=4$. The bounds (A.6) and (A.11) allow us to apply the dominated convergence theorem and conclude that the limit (3.3) gives $\mathcal{B}_{\text{L}}(k)$.
@@
$\mathcal{M}_{\text{NL}}^{\text{E}}\overset{\text{loc}}{\longrightarrow}\mathcal{M}_{\text{L}}^{\text{E}}$ (3.5)
@@ 4 Local limit in Minkowski spacetime
The next task is to investigate the local limit of the Minkowskian correlation functions. In this section we
  1. 1.
    describe the problems we face;
  2. 2.
    show that, on general grounds, if it exists, the limit cannot contain ghosts, while it can contain purely virtual particles;
  3. 3.
    study the limit explicitly in correlation functions, concentrating on the bubble and triangle diagrams.
The outcome is that the local limit is singular with the left option of (2.6), while it gives a local theory with PVPs with the right option (with some caveats).
4.1 The problem
We first note that it is not sufficient to compose the map (3.5) with the continuation E $\rightarrow$ M, because the second step admits a plurality of outcomes:
@@ 4.1 The problem
The analytic continuation E $\rightarrow$ M is excluded by unitarity, as we show below in detail. Among the other options, a special place is reserved to the “average continuation” [10], which defines purely virtual particles. There is a worse possibility, though: that the local limit in Minkowski spacetime does not even exist. In particular, we cannot expect that the limit of (2.9) is just (3.4) with $k\rightarrow k_{\text{M}}$, which is ill defined.
@@ 4.2 No-ghost, pro-PVP theorem
Assume for the moment that the vertices $V_{1}$ and $V_{2}$ are identically one. Making a reflection $p_{4}\rightarrow-p_{4}$ in (2.9) we infer that $B_{\text{M}}(k_{\text{M}})=B_{\text{M}}(k_{\text{M}}^{\ast})$. Because of the property (A.5) and the overall factor $i$, the conjugate $B_{\text{M}}^{\ast}(k_{\text{M}})$ coincides with $B_{\text{M}}(k_{\text{M}}^{\ast})$. Hence, $B_{\text{M}}(k_{\text{M}})=-[B_{\text{M}}(k_{\text{M}})]^{\ast}$. Given that the real part of $B_{\text{M}}(k_{\text{M}})$ vanishes identically, its local limit must vanish as well, if it exists.
Thus, the local limit of $B_{\text{M}}(k_{\text{M}})$ is purely imaginary. If it contained propagating particles or ghosts, its real part would be nontrivial. On the contrary, it is allowed to contain PVPs, because they give a purely imaginary bubble diagram (check appendix B).
@@
A more general version of the argument is based on the optical theorem, which reads $iT+iT^{\dagger}=TT^{\dagger}$, where $S=1+iT$ is the $S$ matrix and $iT$ collects the loop diagrams in matrix form. Consider the “empty” theory (2.1). We call it this way, because it does not propagate any degree of freedom, differently from (2.4). This means $TT^{\dagger}=0$, hence the matrix $T$ is Hermitian: all the loop diagrams are purely imaginary.
Since Re[iTiT]=0=0 for arbitrary finite $\lambda$, the local limit $\lambda\rightarrow\infty$ can only return a local theory with Re[iTiT]=0=0, that is to say, still an empty theory. No physical particles or ghosts can be generated by the limit. Purely virtual particles are allowed, precisely because they satisfy Re[iTiT]=0=0.
@@
$\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}=\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}+\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\hskip 0.72229pt.$ (4.3)
@@
$0=\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}\right]=\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}\right]+\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\right].$ (4.4)
Assume, ad absurdum, that the local limit adds degrees of freedom. Then the optical theorem $iT^{\dagger}=TT^{\dagger}=-2\text{Re}[iTiT$], applied to the limit itself, tells us that $TT^{\dagger}$ is nonzero. For example, in the case of the ordinary bubble diagram with circulating physical particles, formula (B.1) gives
$\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}\right]=-\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{\pi}{4\omega_{\mathbf{p}}\omega_{\mathbf{p}-\mathbf{k}}}\left[\delta(k^{0}+\omega_{\mathbf{p}-\mathbf{k}}+\omega_{\mathbf{p}})+\delta(k^{0}-\omega_{\mathbf{p}-\mathbf{k}}-\omega_{\mathbf{p}})\right].$
@@
$\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}\right]\neq 0,\qquad\frac{\partial}{\partial\lambda}\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{loc}}\right]=0,$
@@
$\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}=\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{PVP}}+\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}},$ (4.5)
and recalling that the bubble diagram (B.2) with circulating PVPs is purely imaginary, the real part of the rest vanishes,
$0=\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}\right]=\text{Re}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\right],$
@@
$\text{Im}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{nloc}}\right]=\text{Im}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{PVP}}\right]+\text{Im}\left[\hskip 0.72229pt\rangle\hskip-1.79993pt{\bigcirc\hskip-1.79993pt}\langle\hskip 0.72229pt\hskip 0.72229pt_{\text{rest}}\right].$
We stress once again that these results are predicated on the assumption that the local limit exists. It turns out that it does exist with the second choice of (2.6), but it does not with the first choice. In the rest of this section we show these facts in detail.
4.3 Local limit: tree diagrams
The Euclidean and Minkowskian actions $S_{\text{E}}$ and $S_{\text{M}}$ are related by $S_{\text{E}}=-iS_{\text{M}}$ (since $^{-S_{\text{E}}}\rightarrow e$$^{iS_{\text{M}}}$ inside the functional integral). We know that the integrated coordinates and momenta remain Euclidean. Thus, the propagator $G_{\text{nl}}^{\text{M}}(p_{\text{M}})$ of the nonlocal Minkowski theory is (2.7) times i-i:
$G_{\text{nl}}^{\text{M}}(p_{\text{M}})=-\frac{i\tau}{Q(P(p_{\text{M}}))},$ (4.6)
where $p_{\text{M}}=(-ip^{0},\mathbf{p})$.
At the tree level, formulas (A.3) and (A.12) give
@@ 4.3 Local limit: tree diagrams
where $\mathcal{P}$ denotes the Cauchy principal value. The top result illustrates the problem with the left choice of (2.6): since we are assuming that the polynomial $P$ is the one of (2.2), the $\lambda\rightarrow+\infty$ limit does not give an acceptable propagator for a local theory. The second choice for (2.6), on the other hand, gives the answer we expect for a PVP.
A polynomial like (2.3) changes the picture. However, we do not have further results to report on this here.
4.4 Local limit: bubble diagram